# Mismatch Quest: Visual and Textual Feedback for Image-Text Misalignment

### v1

@@ Abstract
While existing image-text alignment models reach high quality binary assessments, they fall short of pinpointing the exact source of misalignment. In this paper, we present a method to provide detailed textual and visual explanation of detected misalignments between text-image pairs. We leverage large language models and visual grounding models to automatically construct a training set that holds plausible misaligned captions for a given image and corresponding textual explanations and visual indicators. We also publish a new human curated test set comprising ground-truth textual and visual misalignment annotations. Empirical results show that fine-tuning vision language models on our training set enables them to articulate misalignments and visually indicate them within images, outperforming strong baselines both on the binary alignment classification and the explanation generation tasks. Our method code and human curated test set are available at: https://mismatch-quest.github.io/
@@ 1 Introduction
Figure 2: Qualitative analysis of out-of-distribution results: Showcasing image-text pairs generated by Stable-Diffusion XL [54], Stable-Diffusion 2.1 [66], Adobe Firefly [1] and Composable Diffusion [43] (credits to Chefer et al. [8]) text-to-image models alongside the corresponding textual and visual feedback as predicted by the PaLI 55B model finetuned on TV-Feedback.
@@
To evaluate our alignment model, we construct and publish SeeTRUE-Feedback, a human-annotated test set. Human annotators provide textual explanations and approve visual bounding boxes to delineate misalignments, derived from a mixture of real and synthetic images and texts. Our model outperforms other baselines across all metrics: including 10% increase in alignment Accuracy, 20% increase in Entailment w.r.t gold (human-annotated) textual feedback, and a 2-13% increase in F1 for visual feedback. SeeTRUE-Feedback will be publicly available on our project page. We complement our automated metrics with human ratings through an annotation study on Amazon Mechanical Turk [18], where our model outperforms the competing models by more than 100% improvements on all metrics. Our model also shows strong generalization capabilities with out-of-distribution images and prompts from various advanced T2I models such as Stable Diffusion (SD) v1.0 and v2.1 [66], SD XL [54], Composable Diffusion [43], and Adobe Firefly [1]. Finally, our ablation studies verify the advantage of our multitask training, a single model generating both misalignment labels and feedback for different prompts, compared to training individual models for each task, as well as the effectiveness of our training set filtering strategy.
In sum, our contributions are: (a) a feedback-centric data generation method (ConGen-Feedback); (b) a comprehensive training set (TV-Feedback); (c) a human-annotated evaluation set (SeeTRUE-Feedback), which we also make publicly available; (d) trained models that surpass strong baselines.
@@ 3 Textual and Visual Feedback
Images & Texts
Synthetic & Synthetic
Synthetic & Synthetic
Natural & Natural
Natural & Natural
Natural & Synthetic
Natural & Synthetic
# Instances
1,982,362
56,392
418,653
37,327
577,717
19,825
Image
Positive Caption
A cartoon of a person dressed as a joker with a green coat and a blue tie against a gray background.
A close up of a glass of blue liquid on a table with a gray wall behind it .
A kitchen with cabinets, a stove, microwave and refrigerator.
Two men in Germany jumping over a rail at the same time without shirts.
A duck with a yellow beak is swimming in water.
A bed and a table with a lamp on it are in a room with a window and a view of trees.
Negative Caption
A cartoon of a person dressed as a clown with a green coat and a blue tie against a gray background
A close up of a glass of red liquid on a table with a gray wall behind it.
A kitchen with cabinets, a stove, microwave and a toaster.
Two men in Germany jumping under a rail at the same time without shirts.
A duck with a yellow beak is flying in the air.
A bed and a table with a lamp on it are in a room with a window and a view of a lake.
Misalignment Type
Object
Attribute
Object
Relation
Action
Object
Feedback
The person is dressed as a joker, not a clown
The liquid is blue, not red
The kitchen is missing a toaster, but has a refrigerator.
The men are jumping over a rail, not under it
The duck is swimming, not flying
The room has a view of trees, not a lake
Misalignment in Text
clown
red liquid
toaster
jumping under a rail
duck flying
a view of a lake
Visual Misalignment Detection
[2, 3, 996, 995] joker
[380, 308, 944, 666] blue liquid
[193, 327, 347, 553] refrigerator
[277, 26, 664, 477] two men and [608, 3, 729, 998] a rail
[339, 245, 581, 834] duck swimming
[409, 727, 559, 930] trees
@@ 4.1 Collecting Positive Image-Text Pairs
We compile a set of over a million positive image-text pairs, consisting of synthetic and natural images. Approximately 65% of our examples consist of synthetic images, which were generated by a variety of T2I models from PickaPic [33] and ImageReward [74]. For these images, we employ the PaLI 2 [13] model to predict captions that are aligned with the image.
@@
Finally, we take localized narratives [55], captions offering a detailed point-of-view from the annotators) from ADE20k [84, 85] and OpenImages [34] and transform them into more conventional positive captions. To this end, we apply PaLM 2 [3] with a few-shot prompt (see Appendix B) that rewrites the narratives into standardized captions.
@@ 4.2 LLM Generation of Misaligned Image-Text Pairs and Feedback
2) Generate Misalignment and Textual Feedback. Per chosen misalignment candidate, we instruct PaLM 2 [3] API with few-shot prompts to automatically generate: (a) a contradiction caption that introduces the target misalignment; (b) a detailed explanation of the contradiction; (c) a misalignment cue that pinpoints the contradictory element in the caption; and (d) a label for the visual bounding box to be placed on the image. Our instructions and few-shot prompts are presented in Section B.1.
3) Validate the Generation. Some LLM generations may be inaccurate. To increase the quality of the outputs, we filter out examples based on entailment validation as follows. Textual Entailment [19] models classify whether a hypothesis text is entailed by a premise text. We view this relationship as indicating the degree of semantic alignment. We use an entailment model by Honovich et al. [29] to assess the misalignment between our generated contradicting captions (hypothesis) and the original captions (premise), as well as the alignment between feedback (hypothesis) and caption (premise), as illustrated in Sec. B.2. Only valid contradictions and textual feedback, indicated by low and high entailment scores respectively, are retained.
@@ 5 SeeTRUE-Feedback Benchmark
Model / Split Test Val Test Val Test Val Test
PaLI 5B 0.179 0.216 0.226 0.455 0.471 0.478 0.514
InstructBLIP (FlanT5XL) [21] 0.406 0.387 0.555 0.502 0.481 0.392 0.736
mPLUG-Owl (LLaMa-7B-ft) [79] 0.627 0.580 0.297 0.347 0.435 0.485 0.504
MiniGPT-v2 (7B-ft) [10] 0.463 0.374 0.560 0.583 0.440 0.426 0.676
LLaVa-1.5 (Vicuna-7b) [41] 0.571 0.483 0.173 0.206 0.434 0.484 0.717
PaLI 5B ft. Multitask 0.718 0.882 0.765 0.925 0.488 0.575 0.752
PaLI 17B ft. Multitask 0.749 0.867 0.785 0.921 0.353 0.395 0.770
PaLI 55B ft. Multitask 0.743 0.871 0.762 0.898 0.490 0.550 0.787
Table 2: Comparative performance of image/text alignment models on the SeeTRUE-Feedback Benchmark. “ft.” stands for fine-tuned on TV-Feedback.
@@ 5.2 Evaluation Metrics
  • Textual Feedback Quality: Using BART NLI [35]222huggingface.co/facebook/bart-large-mnli, we measure feedback quality by treating ground truth as the ’premise’ and model predictions as the ’hypothesis’, extracting an entailment score (0-100) as semantic alignment.
  • Misalignment in Text: Using BART NLI, we evaluate shorter textual misalignments, determining the model’s precision in identifying text segments incongruent with the image.
@@
We note that Image-Text Alignment is applied to all 8,100 instances from the SeeTRUE dataset. The other metrics are computed on SeeTRUE-Feedback, containing only misaligned pairs.
@@ 6.1 Models and Baselines
Our experiments span multiple leading vision-language models, examined in both zero-shot and fine-tuned scenarios: MiniGPT-v2 (7B-ft) [9], LLaVa-1.5 (Vicuna-7b [40]), InstructBLIP (FlanT5XL[20], mPLUG-Owl (LLaMa-7B-ft) [80], PaLI Series (5B, 17B, 55B) [13, 12]3337B, 5B, 17B, 55B stands for the numbers of model parameters..
@@
Our work uniquely offers an end-to-end assessment of both textual and visual misalignment. To evaluate baseline models for visual misalignment, we adopt a two-step approach. First, we ask for a textual misalignment description. Then, we employ the GroundingDINO grounding model to extract bounding-box information, since the baseline models do not output a bounding-box.
For the supervised experiments, we fine-tuned PaLI models with the visual question answering task using specific questions (additional fine-tuning details are in Appendix Appendix D). The fine-tuning tasks encompass:
@@
    Textual and Visual Feedback: We use a query for combined feedback: “Describe the misalignments between the image and the text: <text>. The expected response format is ‘<feedback> |||| <misalignment in text> |||| <misalignment in image (bounding-box)>’, aiming to extract detailed feedback and specific misalignment indicators in a single model interaction.
@@ 6.2 Main Results
Figure 5: Qualitative comparison of model outputs on two examples from SeeTRUE-Feedback. The PaLI 55B model, fine-tuned on TV-Feedback, effectively identifies a distinct misalignment related to the tennis player’s action and the relative position between the teddy bear and laptop, demonstrating its refined feedback ability.
@@
Overall, the PaLI models fine tuned on TV-Feedback outperform the baselines on all metrics. For example, Non-PaLI models achieved Feedback NLI scores from 0.406 to 0.627, while PaLI models reached 0.718 to 0.749. The largest, 55B PaLI model achieved the highest performance on the binary alignment classification task. Surprisingly, it underperformed the smaller PaLI models on most feedback generation tasks. Specifically, the smaller PaLI 5B is best performing on the in-distribution testset, but less so on the out-of-distribution examples. The PaLI models gap over the baselines is very large on the textual feedback tasks, but less so on the bounding box task. In future work, we plan to improve the multitasking efficiency of the fine-tuned models.
@@ 6.3 Human Ratings and Auto-Metrics Correlations
Table 3 displays results, highlighting PaLI 55B with top scores in feedback accuracy (75.7%), textual misalignment (80.1%), and visual misalignment detection (63.5%), showcasing superior alignment with human judgments
To evaluate our auto-evaluation metrics, we compared our automated evaluation metrics against 1,750 human ratings. Textual metrics included BART NLI [35], BLEU-4 [50], ROUGE-L [38], METEOR [4], CIDEr [70], BERTScore [83], and TRUE NLI [30], while visual metrics comprised AP, IoU, Precision, Recall, and F1-Score at 0.75 threshold. Figure 6 presents these correlations, correlating human ratings with mean auto-metric scores, identifying BART NLI and F1-Score@0.75 as the most correlated textual and visual metrics, respectively. The analysis highlights a robust correlation between our automated metrics and human ratings, confirming the relevance and reliability of our automatic evaluation measures for our task.
@@ 6.4 Out-of-distribution Generalization
We evaluate our model’s generalization capabilities on 100 ‘in-the-wild’ Text-to-Image (T2I) generations from academic papers [8, 62, 67] and Reddit, created using models like Adobe Firefly [1], Composable Diffusion [43], and Stable Diffusion versions 1.0 and 2.1. Figure 2 shows a selection of these results, with more available Appendix A.
We employed the fine-tuned PaLI 55B model on TV-Feedback to predict textual and visual feedback, and these results were rated by three human annotators following our benchmark protocol (Section 5 and Figure 4).
@@
Figure 6: Correlation analysis between human ratings and automated metrics for feedback evaluation. Subfigures (a) and (b) show textual feedback correlations using metrics like BART NLI and BERTScore, while subfigure (c) displays visual feedback correlations with metrics such as IoU and F1-Score. The X-axis represents annotator agreement, and the Y-axis shows the mean metric scores, identifying the most correlated metrics.

### v2

@@ Abstract
While existing image-text alignment models reach high quality binary assessments, they fall short of pinpointing the exact source of misalignment. In this paper, we present a method to provide detailed textual and visual explanation of detected misalignments between text-image pairs. We leverage large language models and visual grounding models to automatically construct a training set that holds plausible misaligned captions for a given image and corresponding textual explanations and visual indicators. We also publish a new human curated test set comprising ground-truth textual and visual misalignment annotations. Empirical results show that fine-tuning vision language models on our training set enables them to articulate misalignments and visually indicate them within images, outperforming strong baselines both on the binary alignment classification and the explanation generation tasks. Our code and human curated test set are available at: https://github.com/MismatchQuest/MismatchQuest.
@@ 1 Introduction
Figure 2: Qualitative analysis of out-of-distribution results: Showcasing image-text pairs generated by Stable-Diffusion XL [54], Stable-Diffusion 2.1 [66], Adobe Firefly [2] and Composable Diffusion [43] (credits to [8]) text-to-image models alongside the corresponding textual and visual feedback as predicted by the PaLI-X model finetuned on TV-Feedback
To evaluate our alignment model, we construct and publish SeeTRUE-Feedback, a human-annotated test set. Human annotators provide textual explanations and approve visual bounding boxes to delineate misalignments, derived from a mixture of real and synthetic images and texts. Our model outperforms other baselines across all metrics: including 10% increase in alignment Accuracy, 20% increase in Entailment w.r.t gold (human-annotated) textual feedback, and a 2-13% increase in F1 for visual feedback. SeeTRUE-Feedback will be publicly available on our project page. We complement our automated metrics with human ratings through an annotation study on Amazon Mechanical Turk [18], where our model outperforms the competing models by more than 100% improvements on all metrics. Our model also shows strong generalization capabilities with out-of-distribution images and prompts from various advanced T2I models such as Stable Diffusion (SD) v2.1 [66], SD XL [54], Composable Diffusion [43], and Adobe Firefly [2]. Finally, our ablation studies verify the advantage of our multitask training, a single model generating both misalignment labels and feedback for different prompts, compared to training individual models for each task, as well as the effectiveness of our training set filtering strategy. We aim to encourage future works based on the presented methodology for various cross-domain applications, such as enhancing text-to-image processes by providing a feedback signal. Furthermore, it can be utilized to identify and correct incorrect annotations in text-image pairs datasets and refine image captioning models by detecting erroneous captions. In sum, our contributions are: (a) a feedback-centric data generation method (ConGen-Feedback); (b) a comprehensive training set (TV-Feedback); (c) a human-annotated evaluation set (SeeTRUE-Feedback), which we also make publicly available; (d) trained models that surpass strong baselines.
@@ 3 Textual and Visual Feedback
Images & Texts Synthetic & Synthetic Synthetic & Synthetic Natural & Natural Natural & Natural Natural & Synthetic Natural & Synthetic
# Instances 1,982,362 56,392 418,653 37,327 577,717 19,825
Image [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Positive Caption A cartoon of a person dressed as a joker with a green coat and a blue tie against a gray background. A close up of a glass of blue liquid on a table with a gray wall behind it . A kitchen with cabinets, a stove, microwave and refrigerator. Two men in Germany jumping over a rail at the same time without shirts. A duck with a yellow beak is swimming in water. A bed and a table with a lamp on it are in a room with a window and a view of trees.
Negative Caption A cartoon of a person dressed as a clown with a green coat and a blue tie against a gray background A close up of a glass of red liquid on a table with a gray wall behind it. A kitchen with cabinets, a stove, microwave and a toaster. Two men in Germany jumping under a rail at the same time without shirts. A duck with a yellow beak is flying in the air. A bed and a table with a lamp on it are in a room with a window and a view of a lake.
Misalignment Type Object Attribute Object Relation Action Object
Feedback The person is dressed as a joker, not a clown The liquid is blue, not red The kitchen is missing a toaster, but has a refrigerator. The men are jumping over a rail, not under it The duck is swimming, not flying The room has a view of trees, not a lake
Misalignment in Text clown red liquid toaster jumping under a rail duck flying a view of a lake
Visual Misalignment Detection [2, 3, 996, 995] joker [380, 308, 944, 666] blue liquid [193, 327, 347, 553] refrigerator [277, 26, 664, 477] two men and [608, 3, 729, 998] a rail [339, 245, 581, 834] duck swimming [409, 727, 559, 930] trees
@@ 4.1 Collecting Positive Image-Text Pairs
We compile a set of over a million positive image-text pairs, consisting of synthetic and natural images. Approximately 65% of our examples consist of synthetic images, which were generated by a variety of T2I models from PickaPic [33] and ImageReward [74]. For these images, we employ the PaLI [13] model to predict captions that are aligned with the image.
@@
Finally, we take localized narratives [55], captions offering a detailed point-of-view from the annotators) from ADE20k [85, 86] and OpenImages [34] and transform them into more conventional positive captions. To this end, we apply PaLM 2 [4] with a few-shot prompt (see Appendix 0.B) that rewrites the narratives into standardized captions.
@@ 4.2 LLM Generation of Misaligned Image-Text Pairs and Feedback
2) Generate Misalignment and Textual Feedback. Per chosen misalignment candidate, we instruct PaLM 2 [4] API with few-shot prompts to automatically generate: (a) a contradiction caption that introduces the target misalignment; (b) a detailed explanation of the contradiction; (c) a misalignment cue that pinpoints the contradictory element in the caption; and (d) a label for the visual bounding box to be placed on the image. Our instructions and few-shot prompts are presented in Section 0.B.1 .
3) Validate the Generation. Some LLM generations may be inaccurate. To increase the quality of the outputs, we filter out examples based on entailment validation as follows. Textual Entailment [19] models classify whether a hypothesis text is entailed by a premise text. We view this relationship as indicating the degree of semantic alignment. We use an entailment model by Honovich et al. [30] to assess the misalignment between our generated contradicting captions (hypothesis) and the original captions (premise), as well as the alignment between feedback (hypothesis) and caption (premise), as illustrated in Sec. 0.B.2 . Only valid contradictions and textual feedback, indicated by low and high entailment scores respectively, are retained.
@@ 5.2 Evaluation Metrics
  • Textual Feedback Quality: Using BART NLI [35]222huggingface.co/facebook/bart-large-mnli, we measure feedback quality by treating ground truth as the ’premise’ and model predictions as the ’hypothesis’, extracting an entailment score (0-1) as semantic alignment.
  • Misalignment in Text: This metric evaluates the model’s ability to identify specific segments within the text that are not aligned with the corresponding image. Similar to the metric above, we use BART NLI to measure the entailment between the predicted text and the ground truth. The goal is to pinpoint the exact parts of the input text that are sources of misalignment.
@@
We note that Image-Text Alignment is applied to all 8,100 instances from the SeeTRUE dataset. The other metrics are computed on SeeTRUE-Feedback, containing only misaligned pairs. Examples showing our metric calculations can be seen at Figure 5.
Figure 5: Metric results on the SeeTRUE-Feedback, showcasing calculations given the input, ground truth, and PaLI ft. model predictions, with NLI entailment scores calculated with BART NLI. The first row shows a high-scoring success example, while the second highlights a low-scoring failure with incorrect feedback and predictions.
@@ 6.1 Models and Baselines
Our experiments span multiple leading vision-language models, examined in both zero-shot and fine-tuned scenarios: MiniGPT-v2 (7B-ft) [10], LLaVa-1.5 (Vicuna-7b [41]), InstructBLIP (FlanT5XL[21], mPLUG-Owl (LLaMa-7B-ft) [80], PaLI Series [13, 12, 11]. Our methodology introduces a feedback task that, while new, aligns with the capabilities expected of leading VLMs, renowned for their instruction following capabilities. The task’s design mirrors scenarios these models encounter during training, ensuring they’re well-equipped to handle it.
@@
Our work uniquely offers an end-to-end assessment of both textual and visual misalignment. To evaluate baseline models for visual misalignment, we adopt a two-step approach. First, we ask for a textual misalignment description. Then, we employ the GroundingDINO grounding model to extract bounding-box information, since the baseline models do not output a bounding-box. In addition, our fine-tuned model is capable of predicting the feedback along with both textual and visual misalignments in a single inference. To accurately assess our model’s performance alongside the baselines, we report the visual misalignment performance using both the GroundingDINO output and our model’s predictions
For the supervised experiments, we fine-tuned PaLI models with the visual question answering task using specific questions (additional fine-tuning details are in Appendix Appendix 0.E ). The fine-tuning tasks encompass:
@@
    Textual and Visual Feedback: We use a query for combined feedback: “Describe the misalignments between the image and the text: <text>. The expected response format is ‘<feedback> | <misalignment in text> | <misalignment in image (bounding-box)>’ , aiming to extract detailed feedback and specific misalignment indicators in a single model interaction.
Table 2: Comparative performance of image/text alignment models on the SeeTRUE-Feedback Benchmark. “ft.” stands for fine-tuned on TV-Feedback. Legend: ()(*)( ∗ ) marks the performance using PaLI bounding-box detector instead of GroundingDINO used for the baseline models.
Model / Split Test Val Test Val Test Val Test
PaLI-3 [12] 0.18 0.22 0.23 0.46 0.47/0.47* 0.35/0.48* 0.51
InstructBLIP (FlanT5XL) [20] 0.41 0.39 0.56 0.50 0.48 0.39 0.74
mPLUG-Owl (LLaMa-7B-ft) [79] 0.63 0.58 0.30 0.35 0.43 0.48 0.50
MiniGPT-v2 (7B-ft) [9] 0.46 0.37 0.56 0.58 0.44 0.43 0.68
LLaVa-1.5 (Vicuna-7b) [40] 0.57 0.48 0.17 0.21 0.43 0.48 0.72
PaLI-3 ft. Multitask [12] 0.72 0.88 0.76 0.92 0.61/0.49* 0.83/0.57* 0.75
PaLI ft. Multitask [13] 0.75 0.87 0.78 0.92 0.65/0.35* 0.84/0.39* 0.77
PaLI-X ft. Multitask [11] 0.74 0.87 0.76 0.90 0.61/0.49* 0.84/0.55* 0.79
Figure 6: Qualitative comparison of model outputs on two examples from SeeTRUE-Feedback. The PaLI-X model, fine-tuned on TV-Feedback, effectively identifies a distinct misalignment related to the tennis player’s action and the relative position between the teddy bear and laptop, demonstrating its refined feedback ability.
@@ 6.2 Main Results
Overall, the PaLI models fine tuned on TV-Feedback outperform the baselines on all metrics. For example, Non-PaLI models achieved Feedback NLI scores from 0.406 to 0.627, while PaLI models reached 0.718 to 0.749. The largest, PaLI-X [11] model achieved the highest performance on the binary alignment classification task. Surprisingly, it underperformed the smaller PaLI models on most feedback generation tasks. Specifically, the smaller but most recent PaLI-3 model, is best performing on the in-distribution testset, but less so on the out-of-distribution examples. The PaLI models gap over the baselines is very large on the textual feedback tasks, but less so on the bounding box task. In future work, we plan to improve the multitasking efficiency of the fine-tuned models. Figure 5 shows metrics results calculated on SeeTRUE-Feedback examples to give a more clear overview. More details about our metrics and evaluation process are available at Appendix 0.D.
@@ 6.3 Human Ratings and Auto-Metrics Correlations
Table 3 displays results, highlighting PaLI-X with top scores in feedback accuracy (75.7%), textual misalignment (80.1%), and visual misalignment detection (63.5%), showcasing superior alignment with human judgments. In Appendix 0.D, we present the annotators’ agreement chart for each model’s predictions.
We evaluated our auto-evaluation metrics against 1,750 human ratings. Textual metrics included BART NLI [35], BLEU-4 [50], ROUGE-L [38], METEOR [5], CIDEr [70], BERTScore [84], and TRUE NLI [29]. Visual metrics comprised AP, IoU, Precision, Recall, and F1-Score at 0.75 threshold. Figure 7 shows the correlations, identifying BART NLI and F1-Score@0.75 as the most correlated textual and visual metrics, respectively. This analysis confirms the relevance and reliability of our automatic evaluation measures.
@@ 6.4 Out-of-distribution Generalization
We evaluate our model’s generalization capabilities on 100 ‘in-the-wild’ Text-to-Image (T2I) generations from academic papers [8, 61, 67] and Reddit, created using models like Adobe Firefly [2], Composable Diffusion [43], and Stable Diffusion versions 1.0 and 2.1. Figure 2 shows a selection of these results, with more available at Appendix 0.A .
We employed the fine-tuned PaLI-X model on TV-Feedback to predict textual and visual feedback, and these results were rated by three human annotators following our benchmark protocol (Section 5 and Figure 4).
@@
Figure 7: Correlation analysis between human ratings and automated metrics for feedback evaluation. Subfigures (a) and (b) explore textual feedback correlations with metrics like BART NLI and BERTScore, while subfigure (c) illustrates visual feedback correlations with metrics like IoU and F1-Score. The X-axis denotes annotator agreement; the Y-axis shows mean metric scores, identifying most correlated metrics.